How to prove 1 + Cos4x/2 = Cos22x . 

Grade:12

2 Answers

Kartik Sawant
8 years ago
1+cos4x/2= (1+2cos22x -1)/2  because cos2x=2cos2x -1 
 so =2cos22x/2 = cos22x. Hence the proof
    
 
Akhil Chinnu
8 years ago
we know that cos2x=2cos2x -1
replace x by 2x
then cos4x=2cos22x -1
so,(1+2cos22x -1)/2
=(2cos22x)/2
=cos22x